# Optimizing Data Transfer Performance and Energy Efficiency with Deep Reinforcement Learning

### v1

@@ 3.6.2 Analysis of Results in Table 1
Figure 4: Comparison of Various DRL Algorithms trained on offline historical data using F&E Reward (Fairness and Efficiency Reward) and T/E Reward (Throughput-Focused Energy Efficiency Reward), and evaluated in simulation and real-world transfers between Chameleon Cloud (TACC) and UC sites.
Offline Training Energy (KJ) $\downarrow$: This metric captures the total energy consumption over the offline training run. DQN uses the least energy (131 KJ) to converge, whereas DDPG expends the most (294 KJ), aligning with its longer training horizon and heavier GPU load.
Inference Time (ms) $\downarrow$ and Model Inference Energy (J) $\downarrow$: After training, each model’s per-step inference cost is measured. DQN has an inference time of 0.57 ms and about 0.098 J per step. DRQN, a lower mean inference time (0.21 ms), consumes around 0.088 J. PPO, being policy-gradient, has a slower inference time (0.74 ms) but also shows the lowest per-inference energy at 0.088 J.
Energy During Online Tuning (KJ) $\downarrow$: When the agent must adapt to new conditions in a real world setting, this column indicates how much additional energy is required for ongoing learning. PPO displays the smallest overhead (2.6 KJ), while DDPG’s overhead is notably higher (9.18 KJ).
Overall, these findings illustrate each algorithm’s trade-offs in training speed, resource usage, and energy efficiency, providing a basis for selecting the method that best aligns with specific performance and operational requirements.
3.6.3 Performance in Throughput and Energy
Figure 4 compares throughput and energy usage for each DRL algorithm, with the top row showing simulation results and the bottom row showing real-world transfers between TACC and UC sites of Chameleon cloud.
In simulation, DQN achieves moderate throughput (around 6–7.5 Gbps) alongside relatively low energy consumption, suggesting partial generalization from offline data. DDPG often pushes throughput closer to 8 Gbps but at a higher energy cost. PPO provides stable performance near 7 Gbps, striking a balanced trade-off between speed and energy overhead. Its recurrent variant, R_PPO, adapts more effectively to fluctuating conditions, often matching or slightly exceeding PPO’s median throughput while remaining energy-competitive. DRQN likewise benefits from recurrent memory but can incur higher energy usage in some runs, reflecting incomplete generalization.
Under real-world conditions as shown in figure  4(c) and  4(d), these trends observed in simulation persist with additional variance. DDPG continues to seek higher throughput but at a notable energy penalty, while DQN and PPO maintain lower overall throughput and energy use. R_PPO stands out for quickly adjusting concurrency and parallelism, often sustaining consistent throughput in dynamic network conditions. DRQN’s performance improves but still shows episodes of increased energy cost. Overall, each algorithm exhibits a unique throughput–energy balance; where one underperforms, it often indicates that the model has not fully generalized from offline to real network states.
3.6.4 Online Tuning Performance and Final Selection
@@
  • DQN and DDPG initially underperform and display larger fluctuations. Although they improve with more episodes, their cumulative rewards remain lower on average.
@@
3.6.5 Why R_PPO for SPARTA? Training and Inference Cost Analysis
While Table 1 shows that R_PPO has higher training and inference overhead compared to some other algorithms (e.g., DQN or PPO), Figures 4 and 5 highlight its superior transfer performance (higher throughput and/or lower energy during actual data movement). This trade-off can be captured mathematically by examining the total energy cost associated with both training and inference.
Let $E_{\text{train}}(M)$ be the energy (in KJ) required to train model $M$, and $E_{\text{inf}}(M)$ be the energy (in J) consumed per inference step. For a given data-transfer job that requires $S$ inference steps (e.g., $S$ monitoring intervals) and for $T$ total jobs, the overall cost to use model $M$ is:
$\text{TotalCost}(M)\;=\;E_{\text{train}}(M)\;+\;T\times S\times E_{\text{inf}}(M).$
We define the average overhead per transfer as:
@@ 3.6.5 Why R_PPO for SPARTA? Training and Inference Cost Analysis
$\bar{O}(M)\;=\;\frac{\text{TotalCost}(M)}{T}\;=\;\frac{E_{\text{train}}(M)}{T}\;+\;S\times E_{\text{inf}}(M).$
As $T$ grows large (i.e., the algorithm is used repeatedly in production), the training cost $\frac{E_{\text{train}}(M)}{T}$ is amortized over many transfers, and the main factor becomes the per-inference cost $E_{\text{inf}}(M)$. Hence, if R_PPO consistently achieves better throughput or lower transfer energy than other algorithms, its higher upfront training cost will be offset once it is used for a sufficient number of transfers.
3.7 SPARTA Models
We train two SPARTA variants, both employing the R_PPO algorithm but differing in their reward functions as detailed in Section 3.3.3. Specifically:
  • SPARTA-FE:
    This model adopts the Fairness and Efficiency objective (F&E Reward), as defined in Eq. 4. By penalizing excessive packet loss, it promotes balanced bandwidth sharing among competing flows, thereby reducing congestion and indirectly lowering energy waste. As discussed in Section 3.3.3, this reward function encourages higher throughput while mitigating congestion-related inefficiencies.
  • SPARTA-T:
    This variant targets the Throughput-Focused Energy Efficiency objective (T/E Reward) presented in Eq. 5. It explicitly prioritizes maximizing throughput per unit energy consumed by adapting concurrency and parallelism to conserve energy without sacrificing acceptable performance. The energy-based reward formulation in Section 3.3.3 guides the agent to achieve this balance.
Both models share the same state-space design and R_PPO training framework (see Section 3.6), differing only in their reward functions. This flexibility allows network operators to select the SPARTA variant that best aligns with their specific performance and energy-efficiency requirements.
Figure 5: Cumulative reward progression over 500 episodes for agents initially trained on Chameleon Cloud nodes with throughput focused energy objective (T/E) reward signal and subsequently tuned on Cloudlab nodes. The plot illustrates the improvement in cumulative rewards per episode during the tuning process, with each line representing a different agent and shaded areas indicating the variability of the achieved rewards.
4 Evaluation
@@
We assess the performance of SPARTA and compare it with state-of-the-art solutions on multiple testbeds (i.e., CloudLab, Chameleon Cloud, and FABRIC) with different network configurations, measuring throughput, energy consumption, and fairness. In each trial, we send $1000\times 1\text{GB}$ files from a dedicated sender to a receiver. We repeat the entire transfer procedure five times in sequence, collecting performance and resource usage metrics for each run. Our results are shown as distributions aggregated from these five trials.
We compare two static tools (rclone and escp), an online optimizer (Falcon_MP [15]), a historical-data-driven method called 2-phase [11], and our two DRL-based approaches, (SPARTA-T and SPARTA-FE). The 2-phase model typically mines prior historical data to guide parameter tuning in real time, but because we did not have historical datasets in our testbed setup, we initialized it from a midpoint range of concurrency ($cc$) and parallelism ($p$). For both DRL SPARTA agents, we initially set $cc$ and $p$ at a midpoint within the concurrency/parallelism range. If prior knowledge about promising throughput at certain settings is available, these initial values can instead be chosen accordingly. Meanwhile, Falcon_MP starts from a baseline configuration and uses gradient descent to tune concurrency and parallelism, while rclone and escp maintain the same parameters for the duration of each session.
@@
Figure 6: Throughput and energy usage in Chameleon, CloudLab, and FABRIC (1 TB of 1 GB files) using six methods: rclone, escp, Falcon_MP, 2-phase, SPARTA-T and SPARTA-FE. The static tools (rclone, escp) fix $(cc,p)=(4,4)$, while Falcon_MP updates $(cc,p)$ via gradient descent from a baseline. SPARTA-T and SPARTA-FE build on reinforcement learning f. 2-phase typically uses historical logs, but here starts from midpoint settings in the absence of extensive data. Due to a lack of hardware counters in FABRIC, only throughput is reported for that testbed.
4.1 Experimental Setup
@@
Chameleon Cloud: We deploy the sender and receiver on gpu_p100 nodes at TACC, featuring Intel Xeon E5-2670 v3 processors (2 CPUs, 48 threads total), 128 GiB RAM, and two NVIDIA P100 GPUs. Each node also includes 400–1000 GB of local storage and a 10 Gbps NIC (Broadcom NetXtreme II BCM57800). The WAN path between TACC and the University of Chicago sites is shared and can support up to 10 Gbps of throughput.
@@
4.2 Performance Across Testbeds
@@
rclone and escp exhibit relatively high energy usage due to underutilized link capacity resulting in prolonged transfer time. Falcon_MP does better than rclone and escp due to improved throughput and smaller transfer time. SPARTA-T and SPARTA-FE maintains the lowest energy expenditure by directly optimizing for throughput-to-energy ratios. 2-phase remains in a mid-range energy profile, reflecting its partial adaptation without an extensive historical dataset.
4.2.1 CloudLab (Figures 6(c) and 6(d))
@@ 4.2.2 FABRIC (Figure 6(e))
4.3 Fairness in Concurrent Transfers
In many real-world deployments, multiple data flows share the same bottleneck link, making fairness essential for efficient resource allocation. To assess this, we launch concurrent transfers using different optimizers. Figures 7(a), 7(b), and 7(c) show three representative scenarios:
@@ 4.3 Fairness in Concurrent Transfers
4.3.1 Jain’s Fairness Index (JFI)
We quantify fairness using Jain’s Fairness Index (JFI) [56], which measures how evenly bandwidth is distributed among $n$ concurrent flows with throughputs $\{T_{1},T_{2},\dots,T_{n}\}$:
@@ 4.3.1 Jain’s Fairness Index (JFI)
4.3.2 SPARTA-T (T/E Reward) Fairness (Figure 7(a))
We begin by running three concurrent transfers using SPARTA-T, which implements the Throughput-Focused Energy Efficiency (T/E) Reward (see Section 3.3.3). As each flow seeks to maximize its throughput-to-energy ratio, they converge to different throughput levels depending on arrival times. Consequently, the overall Jain’s Fairness Index (JFI) remains moderate because the T/E Reward does not heavily penalize bandwidth imbalances. Moreover, SPARTA-T can show larger fluctuations in JFI when multiple flows share the network, since it adjusts concurrency and parallelism primarily to optimize throughput per energy. Once a flow completes or detects diminishing energy returns, other flows opportunistically claim the freed capacity, yielding strong per-bit energy savings but potentially fluctuate fairness score among simultaneous transfers.
4.3.3 SPARTA-FE (F&E Reward) Fairness (Figure 7(b))
@@
4.3.4 Mixed Algorithms Fairness (Figure 7(c))
@@
5 Conclusion
In this paper, we present SPARTA, a multi-parameter deep reinforcement learning (DRL) framework that automatically tunes concurrency ($cc$) and parallelism ($p$) to boost data transfer performance in high-speed networks. Using real-world state transition data, we design two agent variants: SPARTA-T with a throughput-focused energy efficiency (T/E) reward and SPARTA-FE with a fairness and efficiency (F&E) Reward. These reward signals enable the agents to balance throughput with energy usage while ensuring fair bandwidth sharing. Tested on multiple testbeds, SPARTA consistently outperforms both baseline and state-of-the-art methods, requiring only minimal configuration and showcasing the advantages of adaptive resource allocation.
@@ 5 Conclusion
Although our experiments focus on multiple concurrent transfers on shared links, future work could scale this approach to hundreds or even thousands of transfers across geographically distributed environments. This might involve multi-agent DRL or centralized optimization strategies and also incorporate other transport protocol such as TCP BBR  [23].
In summary, our results show that DRL-based resource allocation can build more sustainable, high-performing data transfer infrastructures that save energy, increase throughput, and maintain fairness.
APPENDIX
Below, we present the key hyperparameters used for each RL algorithm evaluated in our study. These tables provide details on network architectures, learning rates, batch sizes, and other parameters tailored to optimize concurrency ($cc$) and parallelism ($p$) under various performance objectives. All tables in this appendix follow the same format, listing each hyperparameter and its corresponding value.
Table 2: DQN Hyper-parameters
Hyperparameter Value
Model type fully-connected
Activation function ReLU
Buffer size 10000
Model hidden layers [128, 128]
Batch size 32
Train frequency 4
Update interval 1000
Exploration fraction 0.1
Final epsilon 0.02
Max gradient norm 10
Table 3: PPO Hyper-parameters
Hyperparameter Value
Model type fully-connected
Learning rate 0.0003
Activation function ReLU
Number of steps 2048
Batch size 64
Policy hidden layers [128, 128]
Value function hidden layers [128, 128]
Number of epochs 10
Gamma 0.99
GAE lambda 0.95
Clip range 0.2
Entropy coefficient 0.0
Value function coefficient 0.5
Max gradient norm 0.5
Normalize advantage true
Use SDE false
Table 4: DDPG Hyper-parameters
Hyperparameter Value
Model type fully-connected
Learning rate 0.001
Activation function ReLU
Buffer size 1000000
Model hidden layers [400, 300]
Learning starts 100
Batch size 256
Tau 0.005
Gamma 0.99
Train frequency 1
Gradient steps 1
Action noise None
Table 5: Recurrent PPO Hyper-parameters
Hyperparameter Value
Model type Recurrent fully-connected
Learning rate 0.0003
Activation function Tanh
LSTM hidden size 256
Batch size 128
Number of LSTM layers 1
Number of epochs 10
Enable critic LSTM True
Gamma 0.99
GAE lambda 0.95
Clip range 0.2
Value function coefficient 0.5
Entropy coefficient 0.0
Max gradient norm 0.5
Normalize advantage True
Use SDE False
Table 6: DRQN Hyper-parameters
Hyperparameter Value
Model type Recurrent fully-connected
Learning rate 0.001
Activation function ReLU
Buffer size 1000000
Model hidden layers [64, LSTM(64)]
Learning starts 100
Batch size 256
Target update period 4
Gamma 0.99
Tau 0.01
Episodes 500000
Epsilon start 0.1
Epsilon end 0.001
Epsilon decay 0.995
Max episode length 128
Random update True

### v2

@@ III-H1 Offline Training Setup
III-H2 Training and Inference Cost (Table I)
Table I reports three types of overhead:
  • Offline training cost. DQN converges fastest and with the lowest training energy (30 min, 300K steps, 131 KJ). DRQN is slowest (94 min), and DDPG is the most energy-intensive to train (294 KJ), consistent with heavier actor–critic updates and higher GPU utilization.
  • Inference cost. Per-step inference is sub-millisecond for all methods. DRQN has the lowest latency (0.21 ms), while PPO is the slowest (0.74 ms). Inference energy is small across the board (0.088–0.098 J/step).
  • Online tuning cost. When adapting to a new environment, PPO requires the least additional tuning energy (2.6 KJ), whereas DDPG requires the most (9.18 KJ).
Overall, Table I quantifies the compute/energy overhead of each learner and motivates the accuracy–overhead trade-offs discussed next.
III-H3 Transfer Throughput and Energy
Figure 6 compares achieved throughput and transfer energy in both simulation (top) and real transfers (bottom). In simulation, DDPG tends to push higher throughput but with higher energy, while DQN and PPO deliver moderate throughput with lower energy. Recurrent policies (R_PPO, DRQN) are generally more stable under fluctuating conditions because memory helps disambiguate partially observed congestion.
On the real environment, the same trends largely hold with increased variance: DDPG remains aggressive (higher throughput, higher energy), DQN/PPO are more conservative, and R_PPO typically sustains steadier throughput while controlling transfer energy by adapting $(cc,p)$ quickly. Performance gaps between simulation and real runs indicate incomplete generalization from offline data to unseen network regimes, motivating online tuning when deployment conditions change.
@@ III-H3 Transfer Throughput and Energy
Figure 6: Comparison of RL algorithms trained offline and evaluated in simulation and real transfers.
Figure 7: Cumulative reward progression over 500 episodes for agents trained on one environment and tuned on another.
III-H4 Online Tuning Performance and Final Selection
@@
  • DQN and DDPG initially underperform and display larger fluctuations. Although they improve with more episodes, their cumulative rewards remain lower.
@@
III-H5 Why R_PPO for SPARTA? Training and Inference Cost Analysis
Table I shows that R_PPO has higher one-time training cost and slightly higher per-step inference overhead than some baselines (e.g., DQN, PPO). However, Figures 6 and 7 show that R_PPO achieves better transfer-level behavior (higher throughput and/or lower data-plane energy). Since SPARTA is trained once and then used repeatedly, the relevant question is: after how many transfers does R_PPO’s lower transfer energy outweigh its higher training cost?
We quantify this trade-off by accounting for both training energy and per-transfer energy. Let:
  • $E_{\text{train}}(M)$: energy to train model $M$ (J),
  • $E_{\text{inf}}(M)$: inference energy per monitoring interval (MI) step (J/step),
  • $E_{\text{data}}(M)$: data-plane energy per MI step during transfer (J/step),
  • $S$: number of MI steps per transfer,
  • $T$: number of transfers executed using model $M$.
The total energy over $T$ transfers is
$\text{TotalCost}(M)=E_{\text{train}}(M)+T\times S\times\Bigl[E_{\text{inf}}(M)+E_{\text{data}}(M)\Bigr],$ (17)
and the average energy per transfer is
@@ III-H5 Why R_PPO for SPARTA? Training and Inference Cost Analysis
$\scriptsize\bar{O}(M)=\frac{\text{TotalCost}(M)}{T}=\frac{E_{\text{train}}(M)}{T}+S\times\Bigl[E_{\text{inf}}(M)+E_{\text{data}}(M)\Bigr].$ (18)
As $T$ grows, the amortized term $E_{\text{train}}(M)/T$ becomes small, and the dominant cost is the per-transfer data-plane energy.
To find the break-even point where R_PPO becomes more energy-efficient than an alternative model $M^{\prime}$, we solve:
$E_{\text{train}}(\text{R\_PPO})+T\,S\,\Bigl[E_{\text{inf}}(\text{R\_PPO})+E_{\text{data}}(\text{R\_PPO})\Bigr]$ (19)
$<\;E_{\text{train}}(M^{\prime})+T\,S\,\Bigl[E_{\text{inf}}(M^{\prime})+E_{\text{data}}(M^{\prime})\Bigr].$
Rearranging yields:
$\scriptsize T>\frac{E_{\text{train}}(\text{R\_PPO})-E_{\text{train}}(M^{\prime})}{S\Bigl[\bigl(E_{\text{inf}}(M^{\prime})+E_{\text{data}}(M^{\prime})\bigr)-\bigl(E_{\text{inf}}(\text{R\_PPO})+E_{\text{data}}(\text{R\_PPO})\bigr)\Bigr]}.$ (20)
Example (vs. DQN).
Assume $S=1250$ steps per transfer (e.g., transferring 1000 files of 1 GB each). Using Table I and Figure 6(d):
  • R_PPO: $E_{\text{train}}=221\,K\,\text{J}$, $E_{\text{inf}}\approx 0.094\,\text{J/step}$, $E_{\text{data}}\approx 90\,\text{J/step}$.
  • DQN: $E_{\text{train}}=131\,K\,\text{J}$, $E_{\text{inf}}\approx 0.098\,\text{J/step}$, $E_{\text{data}}\approx 110\,\text{J/step}$.
Substituting:
[…]
Thus, after roughly 4 transfers, the training-energy premium of R_PPO is recovered by its lower per-transfer energy. Since SPARTA is intended for repeated use (often hundreds to thousands of transfers), the one-time training overhead is quickly amortized, which justifies selecting R_PPO despite its higher upfront cost.
III-I SPARTA Models
We train two SPARTA variants. Both use the same R_PPO policy architecture and state/action design; they differ only in the reward (Section III-E):
  • SPARTA-FE (Fairness & Efficiency): uses the F&E reward (Eqs. 815) to maximize throughput while penalizing congestion (via packet loss), which promotes fair sharing and reduces wasted work from retransmissions/queueing.
  • SPARTA-T (Throughput-focused Energy): uses the T/E reward (Eqs. 815) to maximize throughput per unit energy, explicitly favoring energy-efficient parameter settings.
This separation lets operators choose the objective that matches the deployment goal (fairness/congestion control vs. energy efficiency) without changing the learning algorithm.
IV Evaluation
@@
We assess the performance of SPARTA and compare it with state-of-the-art solutions on multiple testbeds (e.g., CloudLab, Chameleon Cloud, and FABRIC) with different network configurations, measuring throughput, energy consumption, and fairness. In each trial, we send $1000\times 1\text{GB}$ files from a dedicated sender to a receiver. We repeat the entire transfer procedure five times in sequence, collecting performance and resource usage metrics for each run. Our results are shown as distributions aggregated from these five trials.
@@
Figure 8: Throughput and energy usage in Chameleon, CloudLab, and FABRIC (1 TB of 1 GB files) using six methods: rclone, escp, Falcon_MP, 2-phase, SPARTA-T and SPARTA-FE. The static tools (rclone, escp) fix $(cc,p)=(4,4)$, while Falcon_MP updates $(cc,p)$ via gradient descent from a baseline $(cc,p)=(1,1)$. SPARTA-T and SPARTA-FE build on reinforcement learning . 2-phase typically uses historical logs, but here starts from midpoint settings in the absence of extensive data. Due to a lack of hardware counters in FABRIC, only throughput is reported for that testbed.
We compare two static tools (rclone and escp), an online optimizer (Falcon_MP [7]), a historical-data-driven method called 2-phase [46], and our two DRL-based approaches (SPARTA-T and SPARTA-FE). The 2-phase model typically mines prior historical data to guide parameter tuning in real time, but because we did not have historical datasets in our testbed setup, we initialized it from a midpoint range of concurrency ($cc$) and parallelism ($p$). For both DRL SPARTA agents, we initially set $cc$ and $p$ at a midpoint within the concurrency/parallelism range. If prior knowledge about promising throughput at certain settings is available, these initial values can instead be chosen accordingly. Meanwhile, Falcon_MP starts from a baseline configuration and uses gradient descent to tune concurrency and parallelism, while rclone and escp maintain the same parameters for the duration of each session.
IV-A Experimental Setup
@@
Chameleon Cloud: We deploy the sender and receiver on gpu_p100 nodes at TACC, featuring Intel Xeon E5-2670 v3 processors (2 CPUs, 48 threads total), 128 GB RAM, and two NVIDIA P100 GPUs. Each node also includes 400–1000 GB of local storage and a 10 Gbps NIC (Broadcom NetXtreme II BCM57800). The WAN path between TACC and the University of Chicago sites is shared and can support up to 10 Gbps of throughput.
@@
IV-B Performance Across Testbeds
@@
rclone and escp exhibit relatively high energy usage due to underutilized link capacity resulting in a prolonged transfer time. Falcon_MP does better than rclone and escp due to improved throughput and smaller transfer time. SPARTA-T and SPARTA-FE maintains the lowest energy expenditure by directly optimizing for throughput-to-energy ratios. 2-phase remains in a mid-range energy profile, reflecting its partial adaptation without an extensive historical dataset.
IV-B1 CloudLab (Figures 8(c) and 8(d))
@@ IV-B2 FABRIC (Figure 8(e))
IV-C Fairness in Concurrent Transfers
In many real-world deployments, multiple data flows share the same bottleneck link, making fairness essential for efficient resource allocation. To assess this, we launch concurrent transfers using different optimizers. Figures 9(a), 9(b), and 9(c) show three representative scenarios:
IV-C1 Jain’s Fairness Index (JFI)
We quantify fairness using Jain’s Fairness Index (JFI) [28], which measures how evenly bandwidth is distributed among $n$ concurrent flows with throughputs $\{T_{1},T_{2},\dots,T_{n}\}$: 3w3444444
@@ IV-C1 Jain’s Fairness Index (JFI)
IV-C2 SPARTA-T (T/E Reward) Fairness (Figure 9(a))
We begin by running three concurrent transfers using SPARTA-T, which implements the Throughput-Focused Energy Efficiency (T/E) Reward. As each flow seeks to maximize its throughput-to-energy ratio, they converge to different throughput levels depending on arrival times. Consequently, the overall Jain’s Fairness Index (JFI) remains moderate because the T/E Reward does not heavily penalize bandwidth imbalances. Moreover, SPARTA-T can show larger fluctuations in JFI when multiple flows share the network, since it adjusts concurrency and parallelism primarily to optimize throughput per energy. Once a flow completes or detects diminishing energy returns, other flows opportunistically claim the freed capacity, yielding strong per-bit energy savings but potentially fluctuate fairness score among simultaneous transfers.
IV-C3 SPARTA-FE (F&E Reward) Fairness (Figure 9(b))
@@
IV-C4 Mixed Algorithms Fairness (Figure 9(c))
@@
V Conclusion
In this paper, we present SPARTA, a multi-parameter deep reinforcement learning (DRL) framework that automatically tunes concurrency ($cc$) and parallelism ($p$) to boost data transfer performance in high-speed networks while keeping the energy consumption low. Using real-world state transition data, we design two agent variants: SPARTA-T with a throughput-focused energy efficiency (T/E) reward and SPARTA-FE with a fairness and efficiency (F&E) Reward. These reward signals enable the agents to balance throughput with energy usage while ensuring fair bandwidth sharing. Tested on multiple testbeds, SPARTA consistently outperforms both baseline and state-of-the-art methods, requiring only minimal configuration and showcasing the advantages of adaptive resource allocation.
@@ V Conclusion
Our results show that DRL-based resource allocation can build more sustainable, high-performing data transfer infrastructures that save energy, increase throughput, and maintain fairness. Although our experiments focus on multiple concurrent transfers on shared links, future work could scale this approach to thousands of transfers across geographically distributed environments. This might involve multi-agent DRL or centralized optimization strategies and also incorporate other transport protocols, such as TCP BBR  [14].
Acknowledgments
This project is in part sponsored by the National Science Foundation (NSF) under award number OAC-2313061.